# SPYGLASS. VII-A. The Demographics and Ages of Small Nearby Young Associations

### v1

@@ 3.1.2 Spatial-Photometric Membership Probabilities
SPYGLASS-IV provides photometric youth probabilities, $P_{Age<50Myr}$, in addition to spatial membership probabilities, $P_{spatial}$, which estimate the local membership fraction using the ratio between the number of near-certain young and old stars. $P_{spatial}$ is available for all stars in the candidate member lists in SPYGLASS-IV, while $P_{Age<50Myr}$ is available for a subset of stars that pass the photometric and astrometric quality cuts used in that publication. SPYGLASS-IV calculates $P_{Age<50Myr}$ assuming the demographics of the field, which is taken to have a constant star formation rate over the last 11.2 Gyr (Binney et al., 2000). This implies that 0.4% of stars are younger than 50 Myr, however associations contain young stars at a much higher rate than the field, and that rate is measured by $P_{spatial}$. We therefore produce spatial-photometric membership probabilities, $P_{sp}$, which readjust the initial youth probability prior for a given value of $P_{spatial}$, following the routine presented in SPYGLASS-V and also employed in SPYGLASS-VI.
Contaminants in CasE necessitate adjustments to our probability assessments, as the field sequence there is elevated slightly above the typical field sequence seen in other populations. This may be caused by the presence of one or more older open clusters that share the parameter space and are younger than the rest of the field. The most probable contaminant population is HSC 1040 (see Hunt & Reffert, 2023), which overlaps with CasE in spatial coordinates and is within 3 mas/yr in proper motion space. To avoid potential contamination from these sources, we apply a manual cut, which we present in Section , and set $P_{sp}$ for all sources below it to zero. Due to the youth of CasE, most stars excluded by this cut are more than a magnitude below the association’s PMS, so the risk of false rejections due to this cut is low.
@@ 3.1.3 Velocity Membership
SPYGLASS-IV selected association members in 5-D space-transverse velocity coordinates, and thus, stars with consistent positions and transverse velocities but inconsistent radial velocities may be marked as candidate members. We therefore use the new and literature radial velocities to assess membership.
To do this, we compute a median transverse velocity in $l$ and $b$ across stars with $P_{sp}>0.5$, and compute median UVW galactic cartesian velocities using stars with $P_{sp}>0.5$ and $\sigma_{RV}<1.5$ km s-1. We use RA/Dec rather than ll/$b$ for LeoE and LeoC due to their proximity to the north galactic pole. we take the difference between the median transverse velocity and the transverse velocity of the most outlying candidate member as the radius within which stars are considered members.
@@ 3.1.4 Velocity False Positives and Negatives
Following SPYGLASS-V, we estimate false positive and false negative rates for velocity membership. These quantities can have a substantial influence on the overall membership probability of candidate members, especially on the main sequence, where field contamination is most common.
@@
We compute the false positive rate by producing a histogram of the RVs with $\sigma_{RV}<5$ km s-1, masking bins in the range occupied by association members as per our RV selection, and fitting a Gaussian to the resulting distribution.
, we instead compute an RV, mean, and standard deviation among probable non-members with $P_{sp}<0.5$, or $P_{sp}<0.2$ if the more lenient cut does not remove most of the spike associated with the population. We then use a Gaussian with those parameters to model the background. The false positive rate is provided by the fraction of field stars, as set by the field model, that would fall in the range of RVs where stars are marked as association members by our RV membership assessment. We show our fits to the field RV distributions in Figure 1. The values of this false positive rate range from 9% in Theia 72 and ScuN to 22% in CasE and SCYA-54.
For the false negative rate, we use the 5% value adopted in SPYGLASS-V. This percentage describes the fraction of association members that are expected to have RVs inconsistent with the population, most likely due to velocities induced by the presence of a companion. The small populations discussed in this paper make it difficult to measure this value independently for each association, however this 5% value is consistent with the result in SPYGLASS-V and with other cases where false negatives have been discussed (e.g., Shkolnik et al., 2017).
@@ 3.1.5 Spectroscopic Youth Indicators
To assess Li membership, we use the EAGLES package, which computes age posteriors for stars of a given Li EW and $T_{eff}$ value. The latter can be acquired from $G_{BP}-G_{RP}$ using a conversion presented in Jeffries et al. (2023), which was shown to produce age values that differ from those of directly-measured values of $T_{eff}$ by only 0.01 dex on average. Li EWs in our sample typically do not account for the presence of the Fe I line at 6707.44 Å , but Soderblom et al. (1993) provides a corrective factor as a function of $T_{eff}$ that accounts for this often-unmodeled line. We apply this correction to all Li EWs prior to running EAGLES, except for the GALAH DR4 values, which model the Fe I line (see Wang et al., 2024). EAGLES provides models of the expected Li EW for $3000<T_{eff}<6500$, so our ability to assess youth with EAGLES is limited to that range (Jeffries et al., 2023). Following SPYGLASS-V, we mark stars with EAGLES analytical age posteriors $P_{age<80Myr}<0.005$ as non-members, which is a strict selection that only removes stars inconsistent with membership at ages near the SPYGLASS limit of detectability or with substantial contributions from a Li-free unresolved companion. We additionally require $RUWE<1.2$ for an assessment of non-membership to exempt stars with evidence of an unresolved binary (Lindegren, 2018; Bryson et al., 2020), which reduces false negatives caused by a companion masking the Li line depth. We then mark stars with $\tau_{mean}<50$ Myr as Li-verified members, a cut that reflects the typical age range of SPYGLASS associations. The remaining stars are marked as having ambiguous Li membership. Of the stars marked as Lithium members, 9 of 100 across all groups are marked as RV non-members, which is within the 90% binomial confidence interval of the 5% velocity false negative rate chosen in Section 3.1.4. One of the 3 velocity non-members in TOR-1 later gets reclassified as a velocity member of a dynamically distinct subgroup in Section 3.2.1, bringing these values into even closer agreement.
@@ 3.1.6 Final Membership Probabilities
Finally, we combine the results across all membership indicators to produce aggregate membership probabilities for all stars in the sample. First, we adjust the spatial-photometric membership probabilities using the velocity membership assessment. Velocity is not a conclusive membership indicator due to the non-zero false positive and negative rates discussed in Section 3.1.4, and final membership rates can therefore be computed using Bayes’ Theorem as the posterior probability $P(mem|V=x)$, where $V$ is the velocity membership assessment, following SPYGLASS-V:
@@
where $P_{fn}$ is the false negative rate (5%). The marginal probability $P(V=1)$ can be written as:
@@
For a given star with a given velocity membership assessment $V$, we evaluate the corresponding formula for its membership probability given its value of $V$, with the result providing its combined membership probability. For stars with inconclusive membership probabilities ($V=0$), the membership probability remains $P_{sp}$. A more detailed motivation for these formulae is provided in SPYGLASS-V.
@@
We show CMDs for all 15 associations covered by this publication in Figure 4, . There we also show the cut used for removing field contaminants in CasE. All of the associations, with the exception of SCYA-79 (see Section 4.1.4), show a clear separation between the stars in the field sequence (near the black isochrone), and the young populations defined by $P_{fin}>0.5$.
@@ 4.1.1 Isochronal Ages
Isochronal ages are the most widely available method of age computation, requiring only absolute photometry, which is available for nearly 2 billion stars in Gaia (e.g., Soderblom et al., 2014). While disagreements between different isochronal models result in systematic uncertainties, isochronal ages are quite reliable in relative terms, as small differences in the height of the pre-main sequence can reliably indicate age differences of order a few Myr.
@@
To compute isochronal ages, we follow a generalized form of the fitting routine outlined in SPYGLASS-V, designed for use with several different isochronal models. We restrict our stellar fitting to $1.8<G_{BP}-G_{RP}<4$, covering much of the pre-main sequence, and minimize the influence of binaries by removing stars with $RUWE>1.2$, following Bryson et al. (2020). We further restrict the sample to in cases where this additional restriction does not reduce the number of available stars available below 15, providing a maximally pure sample of single stars in well-populated groups where further restriction does not adversely affect sample size. We also remove stars with mass $M<0.2M_{\odot}$ (see Sec. 4.2.1). Stars in our samples below this cutoff rarely follow the same isochrone as more massive members, and the offset differs by model. Their inclusion results in age solutions that grow older with the number of stars in this mass range, which correlates with distance. Wang et al. (2025) acknowledged this pattern and corrected the PARSEC isochrones to remove it, although here we simply remove these low-mass stars for distance-unbiased results across all models.
@@
For each population, we take 10000 sub-sampled samples of the stars that survive the cuts above, selecting half of those stars at random to a minimum sample size of 5 with selection probability weighted by $P_{spatial}$. We then apply least-squares fitting with 2-$\sigma$ outlier removal to each sample, selecting a best fit isochrone model based on the $G_{BP}-G_{RP}$ and $M_{G}$ absolute photometry of the selected stars, corrected for Gaia geometric distance (Bailer-Jones et al., 2021) and extinction from Lallement et al. (2019). This routine provides a reliable uncertainty measurement while suppressing contamination from any remaining binaries or stars with excess emission from accretion (e.g., Muzerolle et al., 1998). The final age solution and uncertainty are provided by the 2-$\sigma$ clipped mean and standard deviation, respectively, across all samples. We apply this fitting routine to all populations and subgroups included in this paper, using the BHAC15 (Baraffe et al., 2015), DSEP-Magnetic (Feiden, 2016), and PARSEC (Chen et al., 2015) isochronal models. The PARSEC and BHAC15 isochrones are chosen to cover sets of isochrones that produce similar age solutions in Herczeg & Hillenbrand (2015), while the DSEP-Magnetic isochrones are chosen due to their success in reproducing young association ages computed using other methods, with those results accomplished by introducing strong magnetic fields to the model (e.g., Kerr et al., 2022b).
@@ 4.1.2 Dynamical Ages
Dynamical ages are measured by tracing a population back in time through the galactic potential and measuring the time of most compact configuration, which marks the start of stellar dispersal after formation. Couture et al. (2023) reviews the strengths and weaknesses of various dynamical age calculation methods, highlighting the importance of a pure sample of members and a reliable size metric. They note that the presence of observational uncertainties increases the velocity dispersion of a population, resulting in an overestimation of the rate of stellar dispersal. This produces an offset that, if uncorrected, results in age underestimation. We therefore compute dynamical ages that include a correction for this bias.
@@
We first correct our radial velocities for gravitational redshift, which is caused by general relativistic effects in the gravitational field of a star, and convective blueshift, which is caused by brightness differences of rising and falling convective cells producing line asymmetries that slightly blueshift the spectrum. As described in Couture et al. (2023), both effects change the velocity by a few tenths of a km s-1, usually producing a net redshift. We correct for these effects for each star using the relations in Couture et al. (2023), which provide the expected shifts as a function of spectral class. We determine the spectral class for each star by first interpolating the $T_{eff}$ corresponding to its stellar mass for the parent association’s best fit PARSEC isochrone. We then convert from $T_{eff}$ to spectral class using the relationship from (Herczeg & Hillenbrand, 2014), which covers F5 and later YSOs, and use the (Pecaut & Mamajek, 2013) relation for earlier type stars, which evolve quickly onto the main sequence. We do not apply a convective blueshift correction to APOGEE data, as this effect is most prominent at short wavelengths and often negligible in the near-infrared (Meunier et al., 2017). We use these corrected RVs for dynamical traceback.
@@
To ensure that our sample contains only stars with reliable velocities, we restrict it to include only high-probability ($P_{fin}>0.8$) velocity-verified members without evidence for binarity (resolved or unresolved as indicated with $RUWE>1.2$). We only include stars with $\sigma_{RV}<1$ km s-1, and exclude stars with only Gaia RVs, which have been shown to host systematic uncertainties at the 1 km s-1 level among young stars (Kounkel et al., 2023). We use the median length of all branches connecting stars as our distance metric, where branches are defined as a line connecting any two members, similar to the median mutual distance metric from SPYGLASS-II. In Couture et al. (2025)’s study of Tuc-Hor, Carina, and Columba, the mean branch length provides the best combination of precision, accuracy, and dynamical age contrast, but we find that our use of the median instead of the mean produces more consistent results given the more limited datasets in our associations.
@@
Prior to computing dynamical ages, we further restrict the sample to eliminate outliers. We first trace stars back through the MWPotential2014 galactic potential using galpy (Bovy, 2015) and assuming the Hogg et al. (2005) solar motion. We trace back 50 Myr for most populations and 80 Myr for the older populations of AndS, SCYA-54, and SCYA-79. The choice of model was shown in Miret-Roig et al. (2020) to have a minimal impact on dynamical age results for populations like those in our sample, so MWPotential2014 is chosen primarily due to its computational efficiency. Following Couture et al. (2023), we eliminate any objects that deviate more than 3-$\sigma$ from the median velocity and position of the association at any point during that traceback history. Finally, we remove stars that reach their minimum average distance from other stars within 5 Myr of the present day for PARSEC ages $\tau>20$Myr, within 1 Myr for populations with $10<\tau<20$Myr, and at 0 Myr for $\tau<10$Myr. This excludes objects with velocities inconsistent with dispersal from a common origin on a timescale similar to their presumed ages. These cuts usually remove less than 20% of candidates, but these restrictions can be quite impactful in associations with fewer reliable RVs.
@@
We then compute bias corrections to account for observational uncertainties in our uncorrected dynamical age results, roughly following the routine outlined in Couture et al. (2023). We first use galpy to trace each star in the dynamical age calculation back to its PARSEC isochronal age, and construct a model of the association at that time. We center the model at the association’s mean X, Y, Z, and corresponding velocities at its PARSEC age, and set the velocity dispersion to the average dispersion across $\Delta v_{T,l}$ and $\Delta v_{T,b}$. The initial size of the association is more difficult to estimate, as molecular clouds vary in size, and traceback uncertainties are too large to resolve most parent cloud sizes. The mean standard deviation across X, Y, and Z at the time of its uncorrected dynamical age serves as a plausible upper limit, but the cloud size is otherwise largely unconstrained. Known clouds have scales ranging from around 1 pc, like Rho Oph and NGC 1333, up to tens of pc, like some filaments in the Taurus complex (e.g., Ladjelate et al., 2020; Krolikowski et al., 2021; Soler et al., 2023). Without a strong reason to favor a small, compact cloud over a larger filamentary structure, we Monte Carlo sample the initial size with a value drawn from a flat distribution between 0.5 pc and the position standard deviation for each population at its uncorrected dynamical age, which ranges from 3-12 pc.
@@
Couture et al. (2023) find that the initial size does not substantially affect the dynamical age bias, but this conclusion assumes a velocity dispersion of 1 km s-1. Many of our associations have $\sigma_{v}<0.5$ km s-1, and we find that in this regime, the size of this bias is sensitive to the initial size of the population. In the case of CMaN, an initial size of 1 pc produces a mean offset of $\sim 2$ Myr, while a 10 pc size produces an offset of 6 Myr. Populations derived from filamentary cloud complexes may therefore have their dynamical ages systematically underestimated.

### v3

@@ 3.1.2 Spatial-Photometric Membership Probabilities
Kerr23 provides photometric youth probabilities, $P_{Age<50Myr}$, in addition to spatial membership probabilities, $P_{spatial}$, which estimate the local membership fraction using the ratio between the number of near-certain young and old stars. $P_{spatial}$ is available for all stars in the candidate member lists in Kerr23, while $P_{Age<50Myr}$ is available for a subset of stars that pass the photometric and astrometric quality cuts used in that publication. Kerr23 calculates $P_{Age<50Myr}$ assuming the demographics of the field, which is taken to have a constant star formation rate over the last 11.2 Gyr (Binney00). This implies that 0.4% of stars are younger than 50 Myr, however associations contain young stars at a much higher rate than the field, and that rate is measured by $P_{spatial}$. We therefore produce spatial-photometric membership probabilities, $P_{sp}$, which readjust the initial youth probability prior for a given value of $P_{spatial}$, following the routine presented in Kerr24 and also employed in Kerr25.
Contaminants in CasE necessitate adjustments to our probability assessments, as the field sequence there is elevated slightly above the typical field sequence seen in other populations. This may be caused by the presence of one or more older open clusters that share the parameter space and are younger than the rest of the field. The most probable contaminant population is HSC 1040 (see Hunt23), which overlaps with CasE in spatial coordinates and is within 3 mas/yr in proper motion space. To avoid potential contamination from these sources, we apply a manual cut, which we present in Section 3.1.6, and set $P_{sp}$ for all sources below it to zero. Due to the youth of CasE, most stars excluded by this cut are more than a magnitude below the association’s PMS, so the risk of false rejections due to this cut is low.
@@ 3.1.3 Velocity Membership
Kerr23 selected association members in 5-D space-transverse velocity coordinates, and thus, stars with consistent positions and transverse velocities but inconsistent radial velocities may be marked as candidate members. We therefore use the new and literature radial velocities to assess membership.
Our RV coverage is often incomplete and contaminated with field stars, so using the distribution of those measurements to set our membership selection risks significant over or under-estimates of the true size of the distribution. We therefore use the 2-D transverse velocity distributions to predict the overall velocity distribution. To do this, we compute a median transverse velocity in $l$ and $b$ across stars with $P_{sp}>0.5$, and compute median UVW galactic cartesian velocities using stars with $P_{sp}>0.5$ and $\sigma_{RV}<1.5$ km s-1. We use RA/Dec rather than ll/$b$ for LeoE and LeoC due to their proximity to the north galactic pole. Assuming that the maximum extent of the population in 3-D space is the same as the clustering-defined maximum extent in 2-D space, we take the difference between the median transverse velocity and the transverse velocity of the most outlying candidate member as the radius within which stars are considered members. This produces a relatively generous membership determination that will produce some false positives, which we account for in Section 3.1.4.
@@ 3.1.4 Velocity False Positives and Negatives
Following Kerr24, we estimate false positive and false negative rates for velocity membership. These quantities can have a substantial influence on the overall membership probability of candidate members, especially on the main sequence, where field contamination is most common.
@@
We compute the false positive rate by producing a histogram of the RVs with $\sigma_{RV}<5$ km s-1, masking bins in the range occupied by association members as per our RV selection, and fitting a Gaussian to the resulting distribution. Some populations in our sample are tenuous enough that a narrow association component is not favored by RV fitting, and in others, modeling the background using something like a KDE struggles to simultaneously model both the dense population and sparse field. Our approach therefore allows us to fit the field alone, without needing to model the much narrower peak associated with the association.
In populations with substructure that makes excluding the association more challenging (e.g., TOR1, see Sec. 3.2.1), or with a field distribution that is either notably non-Gaussian (e.g. CasE) or particularly tenuous (e.g., LeoC), we instead compute an RV, mean, and standard deviation among probable non-members with $P_{sp}<0.5$, or $P_{sp}<0.2$ if the more lenient cut does not remove most of the spike associated with the population. We then use a Gaussian with those parameters to model the background. The false positive rate is provided by the fraction of field stars, as set by the field model, that would fall in the range of RVs where stars are marked as association members by our RV membership assessment. We show our fits to the field RV distributions in Figure 1. The values of this false positive rate range from 9% in Theia 72 and ScuN to 22% in CasE and SCYA-54. Establishing a non-zero false-positive rate is important to allow the removal of otherwise non-credible RV members, although the factor of two variation between these associations has a relatively limited effect, typically translating to a $\sim 10$% change in final membership probability in Section 3.1.6.
For the false negative rate, $P_{fn}$, we use the 5% value adopted in Kerr24. This percentage describes the fraction of association members that are expected to have RVs inconsistent with the population, most likely due to velocities induced by the presence of a companion. The small populations discussed in this paper make it difficult to measure this value independently for each association, however this 5% value is consistent with the result in Kerr24 and with other cases where false negatives have been discussed (e.g., Shkolnik17).
@@ 3.1.5 Spectroscopic Youth Indicators
To assess Li membership, we use the EAGLES package, which computes age posteriors for stars of a given Li EW and $T_{eff}$ value. The latter can be acquired from $G_{BP}-G_{RP}$ using a conversion presented in Jeffries23, which was shown to produce age values that differ from those of directly-measured values of $T_{eff}$ by only 0.01 dex on average. Li EWs in our sample typically do not account for the presence of the Fe I line at 6707.44 Å , but Soderblom93 provides a corrective factor as a function of $T_{eff}$ that accounts for this often-unmodeled line. We apply this correction to all Li EWs prior to running EAGLES, except for the GALAH DR4 values, which model the Fe I line (see Wang24). EAGLES provides models of the expected Li EW for $3000<T_{eff}<6500$, so our ability to assess youth with EAGLES is limited to that range (Jeffries23). Following Kerr24, we mark stars with EAGLES analytical age posteriors $P_{age<80Myr}<0.005$ as non-members, which is a strict selection that only removes stars inconsistent with membership at ages near the SPYGLASS limit of detectability or with substantial contributions from a Li-free unresolved companion. We additionally require $RUWE<1.2$ for an assessment of non-membership to exempt stars with evidence of an unresolved binary (RUWELindegren18; Bryson20), which reduces false negatives caused by a companion masking the Li line depth. We then mark stars with $\tau_{mean}<50$ Myr as Li-verified members, a cut that reflects the typical age range of SPYGLASS associations. The remaining stars are marked as having ambiguous Li membership. Of the stars marked as Lithium members, 9 of 100 across all groups are marked as RV non-members, which is within the 90% binomial confidence interval of the 5% velocity false negative rate chosen in Section 3.1.4. One of the 3 velocity non-members in TOR-1 later gets reclassified as a velocity member of a dynamically distinct subgroup in Section 3.2.1, bringing these values into even closer agreement.
@@ 3.1.6 Final Membership Probabilities
Finally, we combine the results across all membership indicators to produce aggregate membership probabilities for all stars in the sample. First, we adjust the spatial-photometric membership probabilities using the velocity membership assessment. Velocity is not a conclusive membership indicator due to the non-zero false positive and negative rates discussed in Section 3.1.4, and final membership rates can therefore be computed using Bayes’ Theorem as the posterior probability $P(mem|V=x)$, where $V$ is the velocity membership assessment, following Kerr24:
@@
where $P_{fn}$ is the false negative rate (5%, see Section 3.1.4). The marginal probability $P(V=1)$ can be written as:
@@
For a given star with a given velocity membership assessment $V$, we evaluate the corresponding formula for its membership probability given its value of $V$, with the result providing its combined membership probability. For stars with inconclusive membership probabilities ($V=0$), the membership probability remains $P_{sp}$. A more detailed motivation for these formulae is provided in Kerr24.
@@
We show CMDs for all 15 associations covered by this publication in Figure 4, with photometry corrected for BailerJones21 distance and for reddening using the Lallement19 maps. There we also show the cut used for removing field contaminants in CasE. All of the associations, with the exception of SCYA-79 (see Section 4.1.4), show a clear separation between the stars in the field sequence (near the black isochrone), and the young populations defined by $P_{fin}>0.5$.
@@ 4.1.1 Isochronal Ages
Isochronal ages are the most widely available method of age computation, requiring only absolute photometry, which is available for nearly 2 billion stars in Gaia (e.g., Soderblom14). While disagreements between different isochronal models result in systematic uncertainties, isochronal ages are quite reliable in relative terms, as small differences in the height of the pre-main sequence can reliably indicate age differences of order a few Myr.
@@
To compute isochronal ages, we follow a generalized form of the fitting routine outlined in Kerr24, designed for use with several different isochronal models. We restrict our stellar fitting to $1.8<G_{BP}-G_{RP}<4$, covering much of the pre-main sequence, and minimize the influence of binaries by removing stars with $RUWE>1.2$, following Bryson20. We further restrict the sample to $RUWE<1.1$ in cases where this additional restriction does not reduce the number of available stars available below 15, providing a maximally pure sample of single stars in well-populated groups where further restriction does not adversely affect sample size. We also remove stars with mass $M<0.2M_{\odot}$ (see Sec. 4.2.1). Stars in our samples below this cutoff rarely follow the same isochrone as more massive members, and the offset differs by model. Their inclusion results in age solutions that grow older with the number of stars in this mass range, which correlates with distance. Wang25 acknowledged this pattern and corrected the PARSEC isochrones to remove it, although here we simply remove these low-mass stars for distance-unbiased results across all models.
@@
For each population, we take 10000 sub-sampled samples of the stars that survive the cuts above, selecting half of those stars at random to a minimum sample size of 5 with selection probability weighted by $P_{spatial}$. We then apply least-squares fitting with 2-$\sigma$ outlier removal to each sample, selecting a best fit isochrone model based on the $G_{BP}-G_{RP}$ and $M_{G}$ absolute photometry of the selected stars, corrected for Gaia geometric distance (BailerJones21) and extinction from Lallement19. This routine provides a reliable uncertainty measurement while suppressing contamination from any remaining binaries or stars with excess emission from accretion (e.g., Muzerolle98). The final age solution and uncertainty are provided by the 2-$\sigma$ clipped mean and standard deviation, respectively, across all samples. We apply this fitting routine to all populations and subgroups included in this paper, using the BHAC15 (BHAC15), DSEP-Magnetic (Feiden16), and PARSEC (PARSECChen15) isochronal models. The PARSEC and BHAC15 isochrones are chosen to cover sets of isochrones that produce similar age solutions in Herczeg15, while the DSEP-Magnetic isochrones are chosen due to their success in reproducing young association ages computed using other methods, with those results accomplished by introducing strong magnetic fields to the model (e.g., Kerr22a).
@@ 4.1.2 Dynamical Ages
Dynamical ages are measured by tracing a population back in time through the galactic potential and measuring the time of most compact configuration, which marks the start of stellar dispersal after formation. These measurements typically assume minimal self-gravity (see Sec. 4.2.6), and that on the short timescales between the formation of these stars and today, external gravitational perturbations have a limited effect. Couture23 reviews the strengths and weaknesses of various dynamical age calculation methods, highlighting the importance of a pure sample of members and a reliable size metric. They note that the presence of observational uncertainties increases the velocity dispersion of a population, resulting in an overestimation of the rate of stellar dispersal. This produces an offset that, if uncorrected, results in age underestimation. We therefore compute dynamical ages that include a correction for this bias.
@@
We first correct our radial velocities for gravitational redshift, which is caused by general relativistic effects in the gravitational field of a star, and convective blueshift, which is caused by brightness differences of rising and falling convective cells producing line asymmetries that slightly blueshift the spectrum. As described in Couture23, both effects change the velocity by a few tenths of a km s-1, usually producing a net redshift. We correct for these effects for each star using the relations in Couture23, which provide the expected shifts as a function of spectral class. We determine the spectral class for each star by first interpolating the $T_{eff}$ corresponding to its stellar mass for the parent association’s best fit PARSEC isochrone. We then convert from $T_{eff}$ to spectral class using the relationship from (Herczeg14), which covers F5 and later YSOs, and use the (Pecaut13) relation for earlier type stars, which evolve quickly onto the main sequence. We do not apply a convective blueshift correction to APOGEE data, as this effect is most prominent at short wavelengths and often negligible in the near-infrared (Meunier17). We use these corrected RVs for dynamical traceback.
@@
To ensure that our sample contains only stars with reliable velocities, we restrict it to include only high-probability ($P_{fin}>0.8$) velocity-verified members without evidence for binarity (resolved or unresolved as indicated with $RUWE>1.2$). We only include stars with $\sigma_{RV}<1$ km s-1, and exclude stars with only Gaia RVs, which have been shown to host systematic uncertainties at the 1 km s-1 level among young stars (Kounkel23). We use the median length of all branches connecting stars as our distance metric, where branches are defined as a line connecting any two members, similar to the median mutual distance metric from Kerr22a. In Couture25’s study of Tuc-Hor, Carina, and Columba, the mean branch length provides the best combination of precision, accuracy, and dynamical age contrast, but we find that our use of the median instead of the mean produces more consistent results given the more limited datasets in our associations.
@@
Prior to computing dynamical ages, we further restrict the sample to eliminate outliers. We first trace stars back through the MWPotential2014 galactic potential using galpy (Bovy15) and assuming the Hogg05 solar motion. We trace back 50 Myr for most populations and 80 Myr for the older populations of AndS, SCYA-54, and SCYA-79. The choice of model was shown in MiretRoig20 to have a minimal impact on dynamical age results for populations like those in our sample, so MWPotential2014 is chosen primarily due to its computational efficiency. Following Couture23, we eliminate any objects that deviate more than 3-$\sigma$ from the median velocity and position of the association at any point during that traceback history. Finally, we remove stars that reach their minimum average distance from other stars within 5 Myr of the present day for PARSEC ages $\tau>20$Myr, within 1 Myr for populations with $10<\tau<20$Myr, and at 0 Myr for $\tau<10$Myr. This excludes objects with velocities inconsistent with dispersal from a common origin on a timescale similar to their presumed ages. These cuts usually remove less than 20% of candidates, but these restrictions can be quite impactful in associations with fewer reliable RVs.
@@
We then compute bias corrections to account for observational uncertainties in our uncorrected dynamical age results, roughly following the routine outlined in Couture23. We first use galpy to trace each star in the dynamical age calculation back to its PARSEC isochronal age, and construct a model of the association at that time. We center the model at the association’s mean X, Y, Z, and corresponding velocities at its PARSEC age, and set the velocity dispersion to the average dispersion across $\Delta v_{T,l}$ and $\Delta v_{T,b}$. The initial size of the association is more difficult to estimate, as molecular clouds vary in size, and traceback uncertainties are too large to resolve most parent cloud sizes. The mean standard deviation across X, Y, and Z at the time of its uncorrected dynamical age serves as a plausible upper limit, but the cloud size is otherwise largely unconstrained. Known clouds have scales ranging from around 1 pc, like Rho Oph and NGC 1333, up to tens of pc, like some filaments in the Taurus complex (e.g., Ladjelate20; Krolikowski21; Soler23). Without a strong reason to favor a small, compact cloud over a larger filamentary structure, we Monte Carlo sample the initial size with a value drawn from a flat distribution between 0.5 pc and the position standard deviation for each population at its uncorrected dynamical age, which ranges from 3-12 pc.
@@
Couture23 find that the initial size does not substantially affect the dynamical age bias, but this conclusion assumes a velocity dispersion of 1 km s-1. Many of our associations have $\sigma_{v}<0.5$ km s-1, and we find that in this regime, the size of this bias is sensitive to the initial size of the population. In the case of CMaN, an initial size of 1 pc produces a mean offset of $\sim 2$ Myr, while a 10 pc size produces an offset of 6 Myr. Populations derived from filamentary cloud complexes may therefore have their dynamical ages systematically underestimated.